OpenStudy (anonymous):

Solve the equation. 78 = -2 (m+3) + m

OpenStudy (anonymous):

ok distribute the -2 to the m + 3

OpenStudy (anonymous):

first distribute the -2 so 78 = -2m + 6 + m. then combine like terms. so 78 = m + 6. then isolate the variableby subtracting 6 from both sides, thus m = 72

OpenStudy (anonymous):

wouldnt it be -72?

OpenStudy (anonymous):

you would get 78 = -2m - 6 + m

OpenStudy (anonymous):

-2m + m = -m

OpenStudy (anonymous):

-39 = 2m + 3 -42 = 2m m = -21

OpenStudy (anonymous):

A. –28 B. –42 C. –72 D. –84

OpenStudy (anonymous):

78 = -m - 6

OpenStudy (anonymous):

72=-m

OpenStudy (anonymous):

m= - 72

OpenStudy (anonymous):

-72 is the right answer, they must have read your question wrong